OpenStudy (anonymous):

Simplify: (16x5 + 8y2) – 7x5

OpenStudy (unklerhaukus):

\[(16x^5 + 8y^2) – 7x^5\]

Parth (parthkohli):

Easy, just operate on the like terms.

OpenStudy (anonymous):

answer plz.!

OpenStudy (goformit100):

it's correct

OpenStudy (unklerhaukus):

pull off these brackets they are useless \[(16x^5 + 8y^2) – 7x^5\]\[=16x^5 + 8y^2 – 7x^5\] then group like terms \[=(16-7)x^5 + 8y^2\] simplify the brackets

Parth (parthkohli):

To add, you only take off these brackets without changing the inner stuff when there is no negative sign outside that.

Parth (parthkohli):

\( \color{Black}{\Rightarrow -(a + b) = -a - b }\)

OpenStudy (unklerhaukus):

how is that relevant parth/

OpenStudy (anonymous):

still waiting on the answer thou.????

Parth (parthkohli):

Well. If she tackles such questions, then she'll be able to perform them.

OpenStudy (unklerhaukus):

you cant simplify \((16-7)=\cdots\) by your self @Quanda

OpenStudy (anonymous):

A. 24xy – 7x5 B. 9x5 + 8y2 C. 9x + 8y D. 24x5y2 + 7x5 those are the answer choices now just tell me which one.!

OpenStudy (unklerhaukus):

as turns out \(16−7=9\) so \((16−7)=9\) so \((16−7)x^5+8y^2\) \(=⋯x^5+8y^2\)

OpenStudy (anonymous):

yes...@UnkleRhaukus is letting u do the last and easiest step. 16-7 equals wat? and then that number in front of x^5 + 8y^2 is the answer.

OpenStudy (anonymous):

another question.? An 8-pound bag of tomatoes costs $5.60. At the same rate, what would be the cost of 11 pounds of tomatoes?

OpenStudy (anonymous):

$5.60 for 8-pound bag of tomatoes $5.60/8=$0.70 or 70 cents for 1-pound bag of tomatoes since we already know 8 pound bag costs $5.60 we need 3 more pounds (0.7 x 3) = $2.1 total for 11 pound bag of tomatoes is therefore $5.60 + $2.10 = $7.70

OpenStudy (anonymous):

Evaluate the expression: [(7 x 6) + 15]2 + 8

OpenStudy (unklerhaukus):

do the multiplication first

OpenStudy (unklerhaukus):

then add the reaming terms in the square brackets

OpenStudy (unklerhaukus):

square this and add the 8 last

OpenStudy (anonymous):

OMG thanks a lot :D.!

OpenStudy (anonymous):

-2b<-24 Slove the inequality.?